OpenStudy (anonymous):

Connor went to New York for vacation. He spent 3 nights at a hotel and rented a car for 4 days. Jillian stayed at the same hotel, but spent 4 nights and rented a car for 5 days from the same company. If Connor paid $675 and Jillian paid $875, how much did one night at the hotel cost?

OpenStudy (anonymous):

about 175

OpenStudy (♪chibiterasu):

Cost of Hotel = H Cost of Rent = R Days Rented = D Nights Rented= N How much Paid = P So.... P = H x N + R x D Connor's: $675 = H x 3 nights + R x 4 days Jillian's: $875 = H x 4 nights + R x 5 days Solve: ($675 = H x 3 nights + R x 4 days) x (5) = $3,375 = 15 nights x H + 20 days x R ($875 = H x 4 nights + R x 5 days) x (-4) = - $3,500 = -16 nights x H + -20 days x R $3,375 = 15 nights x H + 20 days x R - $3,500 = -16 nights x H + -20 days x R -$125 = -1 nights x H H = $125

OpenStudy (♪chibiterasu):

@lubavasilyuk Please don't give out straight answers.

OpenStudy (anonymous):

...

OpenStudy (♪chibiterasu):

Open STUDY

OpenStudy (♪chibiterasu):

The purpose of this website is to help guide people, not to give them cheat sheets,

OpenStudy (♪chibiterasu):

Have you read the Code of Conduct?

OpenStudy (♪chibiterasu):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer). Please do not copy and paste your problems. If people are going to spend time to help you, you should at least spend the time to ask your question as clearly as possible.

OpenStudy (♪chibiterasu):

Quoted from the CoC

OpenStudy (♪chibiterasu):

I didn't use an aggressive tone, did I? Therefore, I was not mean.

OpenStudy (♪chibiterasu):

17

OpenStudy (♪chibiterasu):

Deleting means nothing, mods and admins can still see your posts.

OpenStudy (anonymous):

3n + 4d = 675 -->(-5) 4n + 5d = 875 --> (4) ------------------- -15n - 20d = -3375 (result of multiplying by -5) 16n + 20d = 3500 (result of multiplying by 4) -------------------add n = 125 One night at the hotel costs $125